# Real-Time Coordination of Integrated Transmission and Distribution Systems: Flexibility Modeling and Distributed NMPC Scheduling

### 01 Feb 2024 v1

@@ IV Case Study
For an illustrative case study, copies of the 118 bus system from the PGLib-OPF [27] are used for the transmission systems and the IEEE 33 bus system for the distribution systems. Four transmission systems are interconnected through multiple connecting tie-lines, as shown in Fig. 1(b), with each connected to 10 distribution systems, roughly representing the 4 TSOs as in Germany. Consequently, the ITD system comprises a total of $1792$ buses. By applying the proposed flexibility model of distribution networks, the resulting MPOPF problems, in the end, will solve dispatch schedules for a system having $472$ buses at the transmission level. Simulations are run for 24 hours with a prediction horizon of $N_{k}=96$ and a time interval of $\Delta t=15$ min. The resulting optimization problems in the transmission level have $187,776$ state variables partitioned into $384$ subproblems. To simulate modern contemporary power system scenarios, we utilize measurement data from the ENTSO-E Transparency Platform [17] dated July 24, 2023 [28]. This day was characterized by heavy rainfall and stood out for relatively significant large prediction mismatches, particularly in solar generation. The dataset comprises day-ahead predictions and actual values for load demand, solar generation, and wind generation, as shown in Fig. 3.
@@
Figure 4: Optimal economic dispatch by isolated approach for 4 TSOs during simulation
Figure 5: Optimal economic dispatch by coordinated approach for 4 TSOs during simulation
Figure 6: State-of-Charge by coordinated approach for 4 TSOs during simulation
@@
In the case study, three distinct operating strategies are introduced and analyzed, i.e., isolated operation, centralized coordination, and distributed coordination. In all these strategies, the flexibility of distribution systems is aggregated into transmission systems as proposed in Sec. II. In the isolated operation mode, transmission systems are operated independently without communication or power sharing among different transmission systems. The results of the economic dispatches per time period are comprehensively visualized in Fig. 4.
Centralized and distributed coordinations employ the system model in (II-B2), treating all systems as a single grid at the transmission level. Unlike isolated operation, the coordinated operating modes facilitate power exchange between transmission systems to minimize overall generation costs, as illustrated in Fig. 5. The difference lies in the optimization approaches. Centralized coordination communicates all private data to a centralized entity and employs a centralized approach to solve the optimization problem. In contrast, distributed coordination solves the optimization problem based on the proposed algorithm in a distributed fashion with limited information exchanged.
@@
A noteworthy instance occurs at $13:00$ where transmission system $T_{1}$ encounters a significant prediction mismatch. In this time period, $T_{1}$ experiences higher actual demand and reduced solar generation, coinciding with lower SOC of ESS s in $T_{1}$, as visualized in Fig. 3 and Fig. 6. In response to this prediction error, power export to other systems is intentionally curtailed as a compensatory measure.
A comparison of economic efficiency is presented in Table I. Notably but not surprisingly, isolated operation stands out with the highest total cost. In contrast, centralized coordination can attain the lowest overall cost. Meanwhile, distributed coordination offers an intriguing alternative, preserving data privacy with limited information exchange while achieving competitive overall costs. Both coordination approaches can find a local minimizer of the system model (II-B2), and optimality gaps in terms of total costs are negligible.
@@
The detailed convergence performance of the proposed distributed approach over $24$ hours is demonstrated in Fig. 7, with the centralized solutions employed as reference solutions. Notably, all the 96 optimization problems in the daily operation exhibit fast convergence in a dozen iterations, within 500 seconds, while both the primal and dual residuals converge to tolerable values. Compared with centralized coordination, the proposed distributed approach showcases remarkable accuracy in terms of controllable power injections and total cost gap over all 96 time periods. These results highlight the scalability and numerical robustness for real-world applications in large-scale ITD systems.
@@ V Conclusion
This paper proposes a distributed framework designed for efficient coordination of ITD systems. Within the framework, the flexibility inherent in distribution systems is aggregated and provided to the respective transmission systems, and the coordination of TSOs is determined by employing an aladin-type distributed nmpc scheme, where the resulting optimization problems are divided across different TSOs and time periods. Numerical simulations conducted on real operational data with relatively large prediction mismatches indicate our approach’s practical applicability in real-world scenarios.

### 20 Feb 2024 v2

@@ IV Case Study
This section presents a comprehensive evaluation of the proposed coordination strategy by examining its performance through operational scenarios on a summer day in Germany, characterized by considerable prediction mismatches due to severe weather conditions.
IV-A Simulation Setting
To model an operational scenario within the German electrical grid, we utilize four 118-bus systems from the PGLib-OPF dataset [36], representing the transmission systems. These are interconnected through multiple tie-lines, reflecting the configuration of the four TSOs in Germany, as depicted in Fig. 1(b). Additionally, each transmission system is connected to 10 distribution systems in a star configuration, employing the IEEE 33-bus system with multiple DERs for these distribution networks. As a result, the ITD system encompasses a total of $1792$ buses with $472$ buses at the transmission level and $1320$ buses at the distribution level.
To capture modern and contemporary power system dynamics under the impact of severe weather, we utilize measurement data from the ENTSO-E Transparency Platform [^2] The data utilized in this paper is available online at the ENTSO-E Transparency Platform: https://transparency.entsoe.eu [24] dated July 24, 2023. As depicted in Fig. 3, the utilized data includes day-ahead predictions (dotted lines) and actual values (solid lines) for load demand, solar generation, and wind generation in each TSO in Germany. This day was marked by adverse weather events, including heavy rainfall, leading to noticeable prediction mismatches, particularly in solar generation. This is visually represented in Fig. 3, highlighting a substantial mismatch during the noon hours.
The simulations cover a $24$-hour period with a prediction horizon with $N_{k}=96$ and a time interval of $\Delta t=15$ min. By aggregating flexibility from DSOs to the transmission level, we significantly reduce the complexity of the optimization problems by not delving into the detailed network topologies but rather by considering the power-energy envelope of the distribution systems at PCC. Consequently, The optimization tasks at the transmission level involve $187,776$ state variables divided into $4$ transmission systems, each across $96$ time periods, resulting in a total of $384$ subproblems.
Note that, in Fig. 3 to Fig. 6, the data are arranged in multiple columns to enable a detailed comparative analysis. Specifically, the first four columns in each figure correspond to data from four distinct control areas, i.e., four TSOs and their respective DSOs, respectively. The final column integrates this data, offering a synthesized overview of these four control areas. This configuration facilitates a straightforward comparison across the spatial decomposition to ensure a structured and clear presentation of the simulation results.
@@ IV-A Simulation Setting
Figure 4: Power generation for optimal economic dispatch by isolated operation mode for 4 TSOs during simulation
Figure 5: Power generation for optimal economic dispatch by coordinated operation mode for 4 TSOs during simulation
Figure 6: State-of-Charge by coordinated operation mode for 4 TSOs during simulation
@@
IV-B Isolated vs. Coordinated Operation Mode
Three distinct operating strategies are explored in the case study: isolated operation, centralized coordination, and distributed coordination. In all these strategies, the flexibility of distribution systems is aggregated to the transmission level as proposed in Sec. II, and the dispatch problems at the transmission level are optimized with a receding horizon. The primary differences between these strategies lie in their operational methodologies and how they address the economic dispatch problems at the transmission level.
In isolated operation mode, each TSO operates in an islanded manner without any communication or power exchange with other transmission systems. The results of the economic dispatches per time period are comprehensively visualized in Fig. 4. The net power generations—calculated as the positive stacked bars minus the negative stacked bars—marginally exceed the actual demands (red lines) over a 24-hour period, across all instances in Fig. 4. This indicates that the balance between supply and demand is maintained, with minimal power line losses.
Contrary to isolated operation, both the centralized and the distributed coordinations utilize the combined system model (II-B2) to facilitate autonomous power exchange (purple bars) between TSOs, aiming to minimize overall generation costs, as depicted in Fig. 5. Fig. 6 demonstrates the state of charge (Soc) of ESS s, highlighting the effective autonomous management in supporting dispatch tasks while adhering to the energy constraints of the ESS s.
@@
A noteworthy instance occurs at $13:00$, highlighted as vertical dotted lines, where transmission system $T_{1}$ encounters a significant prediction mismatch. In this time period, $T_{1}$ experiences higher actual demand and reduced solar generation, as shown in Fig. 3, coinciding with lower SOC of ESS s in $T_{1}$, as shown in Fig. 6. In response to this prediction error, power export to other systems (purple bar) is intentionally curtailed as a compensatory measure, demonstrating the system’s capacity to adapt to unexpected operational dynamics.
IV-C Centralized vs. Distributed Coordination Approaches
The key difference between these two coordination strategies lies in the optimization approaches. Centralized coordination communicates all private data to a centralized entity and employs a centralized algorithm to solve the optimization problem. In contrast, distributed coordination solves the optimization problem based on the proposed algorithm in a distributed fashion with limited information exchanged between TSOs.
@@
Given that both the centralized and the distributed coordination adopt the same system model (II-B2) with $187,776$ state variables divided into 383838384 subproblems, we use centralized solutions as reference solutions to evaluate the effectiveness of the proposed distributed approach in solving the economic dispatch problems at the transmission level. The convergence performance of the proposed distributed approach across $24$ hours is demonstrated in Fig. 7, representing a number of iterations to converge, total computing time for solving one economic dispatch problem, primal and dual residual (III-B), deviations of controllable power injections and optimality gap for each TSOs, expressed as $\left|\frac{f(x_{\ell})-f(x_{\ell}^{*})}{f(x_{\ell}^{*})}\right|$. Notably, all the 96 optimization tasks during the daily operation demonstrate fast convergence in a dozen iterations, under 500 seconds, with both the primal and dual residuals reaching tolerable values. Compared with centralized coordination, the proposed distributed approach showcases remarkable accuracy in terms of controllable power injections and total optimality gap over all 96 time periods. These results highlight the scalability and numerical robustness for real-world applications in large-scale ITD systems.
@@
The economic efficiency comparison among the three operational strategies, as shown in Table I indicates that operating in isolation leads to the highest total costs, whereas centralized coordination results in the lowest. Distributed coordination presents a viable alternative, balancing data privacy and competitive costs, approximately $0.0006\%$ higher than centralized methods. Both coordination strategies effectively find local minimizers of the system model (II-B2), with negligible differences in total costs.
@@ V Conclusion
This paper proposes a novel real-time distributed operational framework for efficient coordination of ITD systems. It employs a flexibility aggregation method at the distribution level, leveraging controllable devices through power-energy envelopes provided by DSOs, thereby avoiding additional computational complexity of economic dispatch problems at the transmission level. Furthermore, the framework’s receding horizon strategy enhances its robustness against prediction mismatches, especially under severe weather conditions, highlighted by a case study of a summer day in Germany. By utilizing real operational data with significant prediction mismatches, this study confirms the framework’s practical relevance and applicability in real-world scenarios. Future work includes further exploring flexibility aggregation methods, utilizing more detailed transmission grid data, and strengthening cyber-physical security.